• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name zhangyi

    Log entry time 01:39:59 on February 10, 2009

    Entry number 260149

    keyword=Feb9 Swing shift summary (late entry)

    16:00 shift start with Moller measurement.
    16:20 Moller fininshed. Vince asked for control access.
    18:30 Control access was finished. Ask for beam back.
    19:21 beam back, start turning beam.
    20:00 fixed the spot++ problem and restart turning beam.
    21:12 Just finished the procedures of beam turning on the run-list, MCC
    found the compton rate was too high. So ask them work on it. Later
    MCC called us saying someone asked a control access again.  (see log
    entry 260115)
    22:00 MCC asked to change target to Optics again, check for beam
    position and raster size check. 
    22:21 Beam come back again, start with Optics target for beam
    checking.
    22:26 move to He3 target, stat with 6x6 raster to check if the rates
    were same. Then go to 2uA 4x4, start taking elastic data.
    22:41 Start the first good run#1203.
    23:59 shift ends with 4 good run:  1203,  1204,  1205 and  1206,
    both of them have 1M event. Run 1207 was started...